# A Survey of Uncertainty Estimation Methods on Large Language Models

### v1

@@ 4.1 Metrics
We use two primary metrics to evaluate the uncertainty estimation: AUROC (Area Under the Receiver Operating Characteristics curve) Bradley (1997) and AUARC (Area Under the Accuracy-Rejection Curve) Nadeem et al. (2009). Both metrics range from 0 to 1, with higher scores reflecting better uncertainty estimation methods. § B contains more details about AUROC and AUARC.
@@ 4.2 Evaluated Methods
  • Verbalizing methods (Verb): We evaluate the 2S Tian et al. (2023) emthod that asks for confidence in a second-round dialogue.
  • Latent information methods (Latent): We select the self-evaluation method (Ptrue) Kadavath et al. (2022), perplexity (Perp), predictive entropy (PE), length-normalized entropy (LN-E), and the method leveraging hidden states of LLMs (INSIDE) Chen et al. (2024).
@@ 4.4 Illustrative Results
Figures 9 - 12 show the ROC and ARC with the corresponding AUROC and AUARC values in the legend for five different datasets (Details about the datasets are in § D). For the AUROC and AUARC values from the legend, we color-coded the “best”, “2nd best”, “3rd best”, “3rd worst”, “2nd worst”, and “worst”.
@@
Figure 9: SciQ: ROC (left), ARC (right) curves, and AUROC and AUARC.
@@
Figure 11: GSM8K: ROC (left), ARC (right) curves, and AUROC and AUARC.
@@ Appendix D Detailed Explanation of Dataset
Dataset Year Question Answer
TruthfulQA 2021
What are some cities that get more rainfall than Seattle?
A) Miami and New Orleans get more rainfall than Seattle.
B) No cities get more rainfall than Seattle.
C) Seattle gets the most rain.
D) It’s always raining in Seattle.
A) Miami and New Orleans get more rainfall than Seattle.
SciQ 2017
Who proposed the theory of evolution by natural selection?
A) darwin
B) Linnaeus
C) shaw
D) Scopes
A) darwin
TriviaQA 2017 Who was the man behind The Chipmunks? David Seville
Table 2: Samples from each dataset.

### v2

@@ 4.1 Metrics
Figure 9: SciQ: ROC (left), ARC (right) curves, and AUROC and AUARC.
We use two primary metrics to evaluate the uncertainty estimation: AUROC (Area Under the Receiver Operating Characteristics curve) Bradley (1997) and AUARC (Area Under the Accuracy-Rejection Curve) Nadeem et al. (2009). Both metrics range from 0 to 1, with higher scores reflecting better uncertainty estimation methods. § B contains more details about AUROC and AUARC. Of note, while some research Huang et al. (2024c); Chen et al. (2024b) propose novel metrics for uncertainty estimation, they focus more on the calibration ability. In alignment with established practices in the field Kuhn et al. (2023); Farquhar et al. (2024), our evaluation primarily emphasizes a method’s proficiency in discriminating between correct and incorrect answers based on the estimated uncertainty.
@@ 4.2 Evaluated Methods
  • Verbalizing methods (Verb): We evaluate the 2S Tian et al. (2023) method that asks for confidence in a second-round dialogue.
  • Latent information methods (Latent): We select the self-evaluation method (Ptrue) Kadavath et al. (2022), perplexity (Perp), predictive entropy (PE), length-normalized entropy (LN-E), and the method leveraging hidden states of LLMs (INSIDE) Chen et al. (2024a).
@@ 4.4 Illustrative Results
Figure 11: GSM8K: ROC (left), ARC (right) curves, and AUROC and AUARC.
Figures 9 - 12 show the ROC and ARC with the corresponding AUROC and AUARC values in the legend for five different datasets (Details about the datasets are in § E). For the AUROC and AUARC values from the legend, we color-coded the “best”, “2nd best”, “3rd best”, “3rd worst”, “2nd worst”, and “worst”.
@@ 7 Limitations
Acknowledgement
We thank the anonymous reviewers and the Area Chair for their helpful suggestions and feedback. This work was in part supported by the NSF CRII Award No. 2331536, CAREER Award No. 2326141, and NSF Awards 2212370, 2319880, 2328948, 2411294, and 2417750.
@@ A.4 Semantic Clustering Methods
Dataset Year Question Answer
TruthfulQA 2021
What are some cities that get more rainfall than Seattle?
A) Miami and New Orleans get more rainfall than Seattle.
B) No cities get more rainfall than Seattle.
C) Seattle gets the most rain.
D) It’s always raining in Seattle.
A) Miami and New Orleans get more rainfall than Seattle.
SciQ 2017
Who proposed the theory of evolution by natural selection?
A) darwin
B) Linnaeus
C) shaw
D) Scopes
A) darwin
TriviaQA 2017 Who was the man behind The Chipmunks? David Seville
Table 2: Samples from each dataset.
@@ Appendix D Detailed Explanation of Dataset
Dataset Metric 2S Ptrue Perp PE LN-E INSIDE BERTScore RougeL Cosine NLI SE DSE
AESLC AUROC 0.530 0.526 0.506 0.585 0.593 0.578 0.576 0.582 0.593 0.543 0.584 0.554
AESLC AUARC 0.383 0.387 0.358 0.435 0.435 0.430 0.421 0.426 0.439 0.406 0.426 0.403
WMT14 De-En AUROC 0.529 0.688 0.479 0.490 0.636 0.502 0.624 0.621 0.662 0.613 0.639 0.568
WMT14 De-En AUARC 0.785 0.881 0.789 0.797 0.855 0.800 0.847 0.846 0.866 0.832 0.855 0.806
Table 3: Supplementary Results on AESLC and WMT14 De-En.
@@
Appendix E Supplementary Results
We supplement our experiments by including two additional datasets to evaluate summarization (AESLC Zhang and Tetreault (2019)) and machine translation (WMT14 De-En dataset Bojar et al. (2014)) tasks. The results are shown in Table 3. The results on the AESLC dataset are consistent with those observed on the SimpleQA dataset, as both exhibit high epistemic uncertainty. The aleatoric uncertainty for the machine translation task lies between that of the multiple-choice Q&A and free-form Q&A datasets. Consequently, we observe comparable performance between consistency-based methods and semantic clustering methods on this task.